# Connecting Federated ADMM to Bayes

### v1

@@ 1 Introduction
We focus on two distinct federated-learning approaches based on the Alternative Direction Methods of Multiplier (ADMM) and Variational Bayes (VB), respectively. The ADMM approach synchronises the global and local models by using constrained optimisation and updates both primal and dual variables simultaneously. This includes methods like FedPD (Zhang et al., 2021), FedADMM (Gong et al., 2022; Wang et al., 2022; Zhou and Li, 2023) and FedDyn (Acar et al., 2021). The VB approach, on the other hand, uses local posterior distributions as messages and multiplies them to compute an accurate estimate of the global posterior (Ashman et al., 2022). This follows the more general Bayesian framework which has a long history in distributed and decentralised computations (Mutambara, 1998; Tresp, 2000; Durrant-Whyte, 2001). Despite decades of work, the two fields remain disconnected and are thought to be fundamentally different: there is no notion of duality in VB, and posterior distributions are entirely absent in ADMM. Little has been done to connect the two at a fundamental level and our goal here is to address this gap.
@@
We use the connection to derive new federated learning algorithms. First, in Sec. 3.3 we extend the isotropic case to learn full-covariances through the federated VB approach. This leads to an ADMM-like approach with an additional dual variable to learn covariances, which are then used as preconditioners for the global update. Second, in Sec. 3.4, building upon functional regularisation approaches from Bayesian literature, we propose to use similar extensions to handle unlabelled data, providing a relationship to federated distillation methods (Seo et al., 2020; Li and Wang, 2019; Lin et al., 2020a; Seo et al., 2020; Wu et al., 2021), as well as work in function-space continual learning (Buzzega et al., 2020; Kirkpatrick et al., 2017; Pan et al., 2020; Titsias et al., 2020). Our empirical results in Sec. 4 show that the new algorithms lead to improved performance (improved convergence) across a range of different models, datasets, numbers of clients, and dataset heterogeneity levels.
We start by introducing federated ADMM and Bayesian federated learning approaches in Sec. 2. We then connect these two approaches together in Sec. 3, where we also derive our improved algorithms. We present empirical results in Sec. 4.
@@ 2 Federated Learning with ADMM and Bayes
The goal of federated learning is to train a global model (parameters denoted by $\bm{w}_{g}$) by using the data distributed across $K$ clients. Ideally, we want to recover the solution that minimises the loss over all the aggregated data, for instance, we may aim for
@@
The Alternating Direction Methods of Multiplier (ADMM) is a popular framework for distributed and federated learning, where the local and global optimisation are performed with constraints that they must converge to the same model. That is, we use constraints $\bm{w}_{k}=\bm{w}_{g}$ enforced through the dual variables, denoted by $\bm{v}_{k}$. The triple $(\bm{w}_{k},\bm{v}_{k},\bm{w}_{g})$ are updated as follows,
@@
$\begin{split}\text{Client updates:}\quad\bm{w}_{k}&\leftarrow\underset{\bm{w}}{\text{argmin}}\;\;\bar{\ell}_{k}(\bm{w})+\bm{v}_{k}^{\top}\bm{w}+\mbox{$\frac{1}{2}$}\alpha\|\bm{w}-\bm{w}_{g}\|^{2},\\ \quad\bm{v}_{k}&\leftarrow\bm{v}_{k}+\alpha(\bm{w}_{k}-\bm{w}_{g}),\quad\text{for all }k\\ \text{Server update:}\quad\bm{w}_{g}&\leftarrow\frac{1}{K}\sum_{k=1}^{K}\left[\bm{w}_{k}+\frac{1}{\alpha}\bm{v}_{k}\right].\end{split}$ (2)
@@
Here, $\alpha>0$ is a scalar and we denote ¯$\bar{\ell}_{k}=\ell_{k}/N_{k}$ where $N_{k}$ is the size of the data at the $k$’th client. Upon convergence, we get $\bm{w}_{g}^{*}=\bm{w}_{k}^{*}$ and $\bm{v}_{k}^{*}=\nabla\bar{\ell}_{k}(\bm{w}_{g}^{*})$ for all $k$.
@@
Such updates are instances of primal-dual algorithms and many variants have been proposed for federated learning (Zhang et al., 2021; Gong et al., 2022; Wang et al., 2022; Zhou and Li, 2023). For example, FedDyn (Acar et al., 2021) is perhaps the best-performing variant which uses a client-specific $\alpha_{k}\propto\alpha/N_{k}$. It also adds an additional hyperparameter $\delta$ (as shown in Eq. 1) through a local weight-decay regulariser added to each local client’s loss. There are also simpler versions (not based on ADMM), for instance, FedProx (Li et al., 2020) where $\bm{v}_{k}$ is not used (or equivalently fixed to 0 in Eq. 2), and even simpler Federated Averaging or FedAvg (McMahan et al., 2016), where the proximal constraint $\bm{w}-\bm{w}_{g}$ is also removed. Such methods do not ensure that $\bm{w}_{k}^{*}=\bm{w}_{g}^{*}$ at convergence. The use of dual-variable is a unique feature of ADMM to synchronise the server and clients.
@@
The Bayesian framework for federated learning takes a very different approach to ADMM. Instead of a point estimate $\bm{w}_{g}^{*}$, the goal for Bayes is to target a global posterior $p_{g}(\bm{w})=p(\bm{w}|\mathcal{D}_{1:K})$ where $\mathcal{D}_{1:K}=(\mathcal{D}_{1},\mathcal{D}_{2},\ldots,\mathcal{D}_{K})$. Instead of loss functions for each $\mathcal{D}_{k}$, we use likelihoods denoted by $p(\mathcal{D}_{k}|\bm{w})$. The solution $\bm{w}_{g}^{*}$ in Eq. 1 can be seen as the mode of the posterior $p_{g}$ whenever there exists likelihood such that $\log p(\mathcal{D}_{k}|\bm{w})=-\ell_{k}(\bm{w})$ and a prior $\log p_{0}(\bm{w})=-\delta\mathcal{R}(\bm{w})$, where both equations need to be true only up to a constant. Therefore, targeting the posterior $p_{g}(\bm{w})$ also recovers $\bm{w}_{g}^{*}$. Often, we compute local posteriors $p_{k}(\bm{w})=p(\bm{w}|\mathcal{D}_{k})$ at clients and combine them at the servers, for instance, Bayesian committee machines (Tresp, 2000) use the following update,
@@
We will derive a connection by using approximate Bayesian approaches, which are required to scale Bayes to large complex models. Specifically, we will use Variational Bayes (VB), although other works focus on Monte Carlo approximations (Al-Shedivat et al., 2021; Deng et al., 2024; Liang et al., 2024). VB finds an approximate posterior, denoted by $q_{g}(\bm{w})$, by minimising the Kullback-Leibler (KL) divergence $\mathbb{D}_{KL}[q_{g}(\bm{w})\,\|\,p_{g}(\bm{w})]$. For exponential-family distributions, the global $q_{g}(\bm{w})$ has a similar form as Eq. 3 but $t_{k}(\bm{w})$ are replaced by likelihood approximations $\hat{t}_{k}(\bm{w})$, also known as “sites”. For instance, when the prior $p_{0}(\bm{w})$ has the same exponential-form as $q_{g}(\bm{w})$, the optimal $q_{g}^{*}(\bm{w})$ is
@@
The sites $\hat{t}_{k}^{*}(\bm{w})$ in VB are parameterised through natural gradients as shown by Khan and Lin (2017, Eq. 11); see further details in Khan et al. (2018, Eq. 18) and Khan and Rue (2021, Sec. 5.4). In this paper, we will show that a distributed estimation of $\hat{t}_{k}^{*}(\bm{w})$ leads to a natural emergence of dual variables and an ADMM-like algorithm. Specifically, we will rely on the Partitioned Variational Inference (PVI) procedure of Ashman et al. (2022) and show that, for a Gaussian $q_{g}(\bm{w})$, we get a line-by-line correspondence to ADMM. We note that the other prominent approach called Expectation Propagation (EP) (Minka, 2001; Guo et al., 2023) also obtains site parameters by optimising the forward KL, $\mathbb{D}_{KL}[p_{g}\,\|\,q_{g}]$, but there is no known connection to the dual variables.
@@ 3.1 Connecting Partitioned Variational Inference (PVI) to ADMM
Partitioned Variational Inference (PVI) (Ashman et al., 2022; Bui et al., 2018) aims to find the best approximation $q_{g}\in\mathcal{Q}$ where $\mathcal{Q}$ is a set of candidate distributions (for example, a set of Gaussian distributions). The goal is to get the VB solution in Eq. 4 but by using an iterative message passing algorithm where local approximations $q_{k}(\bm{w})$ send the sites $\hat{t}_{k}(\bm{w})$ as messages to improve $q_{g}(\bm{w})$. The method update the following triple $(q_{k},\hat{t}_{k},q_{g})$ as shown below,
@@
The update of $q_{g}(\bm{w})$ is exactly the same as Eq. 3 but uses $\hat{t}_{k}(\bm{w})$ which in turn is obtained by using the ratio $q_{k}(\bm{w})/q_{g}(\bm{w})$ in the second line. The $\hat{t}_{k}(\bm{w})$ essentially uses the discrepancy between the global and local distributions. This is then used in the first line to modify the local $q_{k}(\bm{w})$. Similarly to ADMM, at convergence we have $q_{g}^{*}(\bm{w})=q_{k}^{*}(\bm{w})$. The site parameters are related to natural gradients (Khan and Lin, 2017; Khan et al., 2018), therefore we can expect them to be related to dual variables $\bm{v}_{k}^{*}$ in ADMM which estimate gradients. In what follows, we will derive this precisely.
@@
$\begin{split}\text{Client updates:}\quad q_{k}&\leftarrow\arg\min_{q\in\mathcal{Q}}\;\;\mathbb{E}_{q}[\ell_{k}(\bm{w})]+\mathbb{E}_{q}[\log\hat{t}_{k}(\bm{w})]+\mathbb{D}_{KL}[q\,\|\,q_{g}]\\ \log\hat{t}_{k}&\leftarrow\log\hat{t}_{k}+\rho\left(\log q_{k}-\log q_{g}\right)\\ \text{Server update:}\quad\log q_{g}&\leftarrow\log p_{0}+\sum_{k=1}^{K}\log\hat{t}_{k}+\text{const},\end{split}$ (6)
where we have added a damping factor $0<\rho\leq 1$ as it can slow down the rate of change of $\hat{t}_{k}(\bm{w})$, which can be important especially in heterogenous settings (although the theory suggests that $\rho=1$ should be ideal). The three updates have a similar form to the ADMM updates in Eq. 2. The first line uses an additional expectation over $q(\bm{w})$ with the Euclidean distance replaced by the KL divergence. The role of $\hat{t}_{k}(\bm{w})$ appears to be similar to that of the dual term $\bm{v}_{k}^{\top}\bm{w}$ in ADMM. The update of $\hat{t}_{k}(\bm{w})$ in the second line is also very similar, while the update of $q_{g}(\bm{w})$ has some major differences. Next, we will choose the distributions in PVI accordingly to get even closer to the ADMM update.
@@ 3.2 FedLap: A Laplace Version of PVI
Plugging these in Eq. 6 and making a delta approximation (Khan and Rue, 2021, App. C.1), we get FedLap, a Laplace variant of the PVI (we first provide the update, and then its derivation),
@@
The first line is obtained by making the delta approximation, that is, by replacing $\mathbb{E}_{q}[g(\bm{w})]\approx g(\bm{m})$,
@@
    First, instead of using $\alpha$ to update both $\bm{w}_{k}$ and $\bm{v}_{k}$, here we use two separate parameters. The Bayesian update suggests using the prior parameter $\delta$ in the local client update, for instance, we can use weight-decay. For $\rho$, we follow Ashman et al. (2022) and set it to $1/K$.
@@
    Third, the update of $\bm{w}_{g}$ in PVI does not have $\bm{w}_{k}$ in it or a division by $K$. This is due to updates that follow the form in Eq. 3, where we expect different $\hat{t}_{k}$ to automatically weighted. The Bayesian update does not suggest any additional weighting. Despite this difference, FedLap performs as well as the best-performing ADMM method (FedDyn) in our experiments.
Computation and communication cost. FedDyn and FedLap have the same computation and communication cost. The communication cost is $P$ from clients to server (and server to clients), where $P$ is the number of parameters in the model. Each client $k$’s computation cost is dominated by $O(N_{k}PE)$, where $E$ is the number of epochs during optimisation. Computation cost at the global server is negligible, requiring simply adding together vectors of length $P$.
3.3 FedLap-Cov: A New ADMM Variant from VB with Flexible Covariances
Our connection gives a direct way to design a new ADMM variant to improve it: use different candidate set $\mathcal{Q}$, that is, choose different posterior forms for the candidates $q(\bm{w})$. Here, we demonstrate this for multivariate Gaussian $q(\bm{w})=\mbox{${\cal N}$}(\bm{w};\bm{m},\bm{\Sigma})$ where, unlike the previous section, we aim to also estimate the covariance matrix $\bm{\Sigma}$. This extension leads to two dual variables, where the second dual variable acts as a preconditioner similar to the Newton update. We will often write the precision matrix $\bm{S}=\bm{\Sigma}^{-1}$ because it is directly connected to the Hessian (denoted by $\mathbf{H}$) which are more natural for a Newton-like update. In practice, we use a diagonal matrix because it scales better to large models, but our derivation is more general.

### v2

@@ 1 Introduction
We focus on two distinct federated-learning approaches based on the Alternating Direction Method of Multipliers (ADMM) and Variational Bayes (VB), respectively. The ADMM approach synchronises the global and local models by using constrained optimisation and updates both primal and dual variables simultaneously. This includes methods like FedPD (Zhang et al., 2021), FedADMM (Gong et al., 2022; Wang et al., 2022; Zhou and Li, 2023) and FedDyn (Acar et al., 2021). The VB approach, on the other hand, uses local posterior distributions as messages and multiplies them to compute an accurate estimate of the global posterior (Ashman et al., 2022). This follows the more general Bayesian framework which has a long history in distributed and decentralised computations (Mutambara, 1998; Tresp, 2000; Durrant-Whyte, 2001). Despite decades of work, the two fields remain disconnected and are thought to be fundamentally different: there is no notion of duality in VB, and posterior distributions are entirely absent in ADMM. Little has been done to connect the two at a fundamental level and our goal here is to address this gap.
@@
We use this connection to derive new federated learning algorithms. First, we extend the isotropic case to learn full-covariances through the federated VB approach; see Sec. 3.3. This leads to an ADMM-like approach with an additional dual variable to learn covariances, which are then used as preconditioners for the global update. Second, building upon functional regularisation approaches from Bayesian literature, we propose to use similar extensions to handle unlabelled data, providing a relationship to federated distillation methods (Seo et al., 2020; Li and Fedmd, 2019; Lin et al., 2020a; Seo et al., 2020; Wu et al., 2021), as well as work in function-space continual learning (Buzzega et al., 2020; Kirkpatrick et al., 2017; Pan et al., 2020; Titsias et al., 2020); see Sec. 3.4. Our empirical results in Sec. 4 show that the new algorithms lead to improved performance (improved convergence) across a range of different models, datasets, numbers of clients, and dataset heterogeneity levels.
@@ 2 Federated Learning with ADMM and Bayes
We start by introducing federated ADMM and Bayesian federated learning approaches, and then, in Secs. 3 and 4, we will provide new connections along with new algorithms and empirical results. The goal of federated learning is to train a global model (parameters denoted by $\bm{w}_{g}$) by using the data distributed across $K$ clients. Ideally, we want to recover the solution that minimises the loss over all the aggregated data, for instance, we may aim for
@@
The Alternating Direction Method of Multipliers (ADMM) is a popular framework for distributed and federated learning, where the local and global optimisation are performed with constraints that they must converge to the same model. That is, we use constraints $\bm{w}_{k}=\bm{w}_{g}$ enforced through the dual variables, denoted by $\bm{v}_{k}$. The triple $(\bm{w}_{k},\bm{v}_{k},\bm{w}_{g})$ are updated as follows,
@@
$\begin{split}\text{Client updates:}\quad\bm{w}_{k}&\leftarrow\underset{\bm{w}}{\text{argmin}}\;\;\bar{\ell}_{k}(\bm{w})+\bm{v}_{k}^{\top}\bm{w}+\mbox{$\frac{1}{2}$}\alpha\|\bm{w}-\bm{w}_{g}\|^{2},\\ \quad\bm{v}_{k}&\leftarrow\bm{v}_{k}+\alpha(\bm{w}_{k}-\bm{w}_{g}),\quad\text{for all }k,\\ \text{Server update:}\quad\bm{w}_{g}&\leftarrow\frac{1}{K}\sum_{k=1}^{K}\left[\bm{w}_{k}+\frac{1}{\alpha}\bm{v}_{k}\right].\end{split}$ (2)
@@
Here, $\alpha>0$ is a scalar and we denote ¯$\bar{\ell}_{k}=\ell_{k}/N_{k}$ where $N_{k}$ is the size of the data at the $k$’th client. Upon convergence, we get $\bm{w}_{g}^{*}=\bm{w}_{k}^{*}$ and $\bm{v}_{k}^{*}=-\nabla\bar{\ell}_{k}(\bm{w}_{g}^{*})$ for all $k$.
@@
Such updates are instances of primal-dual algorithms and many variants have been proposed for federated learning (Zhang et al., 2021; Gong et al., 2022; Wang et al., 2022; Zhou and Li, 2023). For example, FedDyn (Acar et al., 2021) is perhaps the best-performing variant which uses a client-specific $\alpha_{k}\propto\alpha/N_{k}$. It also adds an additional hyperparameter through a local weight-decay regulariser added to each local client’s loss. There are also simpler versions (not based on ADMM), for instance, FedProx (Li et al., 2020) where $\bm{v}_{k}$ is not used (or equivalently fixed to 0 in Eq. 2), and even simpler Federated Averaging or FedAvg (McMahan et al., 2016), where the proximal constraint $\bm{w}-\bm{w}_{g}$ is also removed. Such methods do not ensure that $\bm{w}_{k}^{*}=\bm{w}_{g}^{*}$ at convergence. The use of dual-variables is a unique feature of ADMM to synchronise the server and clients.
@@
The Bayesian framework for federated learning takes a very different approach to ADMM. Instead of a point estimate $\bm{w}_{g}^{*}$, the goal for Bayes is to target a global posterior $p_{g}(\bm{w})=p(\bm{w}|\mathcal{D}_{1},\mathcal{D}_{2},\ldots,\mathcal{D}_{K})$. Instead of loss functions for each $\mathcal{D}_{k}$, we use likelihoods denoted by $p(\mathcal{D}_{k}|\bm{w})$. The solution $\bm{w}_{g}^{*}$ in Eq. 1 can be seen as the mode of the posterior $p_{g}$ whenever there exists likelihood such that $\log p(\mathcal{D}_{k}|\bm{w})=-\ell_{k}(\bm{w})$ and a prior $\log p_{0}(\bm{w})=-\delta\mathcal{R}(\bm{w})$, where both equations need to be true only up to a constant. Therefore, targeting the posterior $p_{g}(\bm{w})$ also recovers $\bm{w}_{g}^{*}$. Often, we compute local posteriors $p_{k}(\bm{w})=p(\bm{w}|\mathcal{D}_{k})$ at clients and combine them at the servers, for instance, Bayesian committee machines (Tresp, 2000) use the following update,
@@
We will derive a connection by using approximate Bayesian approaches, which are required to scale Bayes to large complex models. Specifically, we will use Variational Bayes (VB), although there also exist work on Monte Carlo methods (Al-Shedivat et al., 2021; Deng et al., 2024; Liang et al., 2024). VB finds an approximate posterior, denoted by $q_{g}(\bm{w})$, by minimising the Kullback-Leibler (KL) divergence $\mathbb{D}_{KL}[q_{g}\,\|\,p_{g}]$. For exponential-family distributions, the global $q_{g}(\bm{w})$ has a similar form to Eq. 3 but $t_{k}(\bm{w})$ are replaced by likelihood approximations $\hat{t}_{k}(\bm{w})$, also known as “sites”. For instance, when the prior $p_{0}(\bm{w})$ has the same exponential-form as $q_{g}(\bm{w})$, the optimal $q_{g}^{*}(\bm{w})$ is
@@
The sites $\hat{t}_{k}^{*}(\bm{w})$ in VB are parameterised through natural gradients as reported by Khan and Lin (2017, Eq. 11); see an exact expression in Khan and Nielsen (2018, Eq. 18) or Khan and Rue (2023, Sec. 5.4). In this paper, we will show that a distributed estimation of $\hat{t}_{k}^{*}(\bm{w})$ leads to a natural emergence of dual variables and an ADMM-like algorithm. Specifically, we will rely on the Partitioned Variational Inference (PVI) procedure of Ashman et al. (2022) and show that, for a Gaussian $q_{g}(\bm{w})$, we get a line-by-line correspondence to ADMM. We note that the other prominent approach called Expectation Propagation (EP) (Minka, 2001; Guo et al., 2023) also obtains site parameters by optimising the forward KL, $\mathbb{D}_{KL}[p_{g}\,\|\,q_{g}]$, but there is no known connection to the dual variables.
@@ 3.1 Connecting Partitioned Variational Inference (PVI) to ADMM
Partitioned Variational Inference (PVI) (Ashman et al., 2022; Bui et al., 2018) aims to find the best approximation $q_{g}\in\mathcal{Q}$ where $\mathcal{Q}$ is a set of candidate distributions (for example, a set of Gaussian distributions). The goal is to get the VB solution in Eq. 4 but by using an iterative message passing algorithm where local approximations $q_{k}(\bm{w})$ send the sites $\hat{t}_{k}(\bm{w})$ as messages to improve $q_{g}(\bm{w})$. The method updates the following triple $(q_{k},\hat{t}_{k},q_{g})$ as shown below,
@@
The update of $q_{g}(\bm{w})$ is exactly the same as Eq. 3 but uses $\hat{t}_{k}(\bm{w})$ which in turn is obtained by using the ratio $q_{k}(\bm{w})/q_{g}(\bm{w})$ in the second line. The $\hat{t}_{k}(\bm{w})$ essentially uses the discrepancy between the global and local distributions. This is then used in the first line to modify the local $q_{k}(\bm{w})$. Similarly to ADMM, at convergence we have $q_{g}^{*}(\bm{w})=q_{k}^{*}(\bm{w})$. The site parameters are related to natural gradients (Khan and Lin, 2017; Khan and Nielsen, 2018), therefore we may expect them to be related to dual variables $\bm{v}_{k}^{*}$ in ADMM which estimate gradients. In what follows, we will derive this precisely.
@@
$\begin{split}\text{Client updates:}\quad q_{k}&\leftarrow\arg\min_{q\in\mathcal{Q}}\;\;\mathbb{E}_{q}[\ell_{k}(\bm{w})]+\mathbb{E}_{q}[\log\hat{t}_{k}]+\mathbb{D}_{KL}[q\,\|\,q_{g}]\\ \log\hat{t}_{k}&\leftarrow\log\hat{t}_{k}+\rho\left(\log q_{k}-\log q_{g}\right)\\ \text{Server update:}\quad\log q_{g}&\leftarrow\log p_{0}+\sum_{k=1}^{K}\log\hat{t}_{k}+\text{const},\end{split}$ (6)
where we have added a damping factor $0<\rho\leq 1$ as it can slow down the rate of change of $\hat{t}_{k}(\bm{w})$, which can be important especially in heterogenous settings (Ashman et al., 2022), although the theory suggests that $\rho=1$ should be ideal. The three updates have a similar form to the ADMM updates in Eq. 2. The first line uses an additional expectation over $q(\bm{w})$ with the Euclidean distance replaced by the KL divergence. The role of $\hat{t}_{k}(\bm{w})$ appears to be similar to that of the dual term $\bm{v}_{k}^{\top}\bm{w}$ in ADMM. The update of $\hat{t}_{k}(\bm{w})$ in the second line is also very similar, while the update of $q_{g}(\bm{w})$ has some major differences. Next, we will choose the distributions in PVI accordingly to get even closer to the ADMM update.
@@ 3.2 FedLap: A Laplace Version of PVI
Plugging these in Eq. 6 and making a delta approximation (Khan and Rue, 2023, App. C.1), we get FedLap, a Laplace variant of the PVI (we first provide the update, and then its derivation),
@@
The first line is obtained by making the delta approximation, that is, $\mathbb{E}_{q\sim\mathcal{N}(\bm{w};\bm{m},\mathbf{I}/\delta)}[g(\bm{w})]\approx g(\bm{m})$,
@@
    First, instead of using $\alpha$ to update both $\bm{w}_{k}$ and $\bm{v}_{k}$, here we use two separate parameters. The Bayesian update suggests using the prior parameter $\delta$ in the local client update, or the overall weight-decay. For $\rho$, we follow Ashman et al. (2022) and set it to $1/K$.
@@
    Third, the update of $\bm{w}_{g}$ in PVI does not have $\bm{w}_{k}$ in it or a division by $K$. This is due to updates that follow the form in Eq. 3, where we expect different $\hat{t}_{k}$ to already be automatically weighted: the Bayesian update does not suggest any additional weighting. Despite this difference, FedLap performs as well as the best-performing ADMM method (FedDyn) in our experiments.
FedDyn and FedLap have the same computation and communication cost. The communication cost is $P$ from clients to server (and server to clients), where $P$ is the number of parameters in the model. Each client $k$’s computation cost is dominated by $O(N_{k}PE)$, where $E$ is the number of epochs during optimisation. Computation cost at the global server is negligible, requiring simply adding together vectors of length $P$.
3.3 FedLap-Cov: A New ADMM Variant with Flexible Covariances
Our connection gives a direct way to design new ADMM variants to improve it: use a different candidate set $\mathcal{Q}$, that is, choose different posterior forms for the candidates $q(\bm{w})$. Here, we demonstrate this for multivariate Gaussian $q(\bm{w})=\mbox{${\cal N}$}(\bm{w};\bm{m},\bm{\Sigma})$ where, unlike the previous section, we aim to also estimate the covariance matrix $\bm{\Sigma}$. This extension leads to two dual variables, where the second dual variable acts as a preconditioner similar to the Newton update. We will often write the precision matrix $\bm{S}=\bm{\Sigma}^{-1}$ because it is directly connected to the Hessian (denoted by $\mathbf{H}$) which is more natural for a Newton-like update. In practice, we use a diagonal matrix because it scales better to large models, but our derivation is more general.